OpenStudy (anonymous):

Hannah has a jar filled with 150 coins made up of quarters and nickels. The total value of the coins is $28.30 How many quarters and how many nickels are in the jar? (Points : 1) 46 quarters; 104 nickels 75 quarters; 75 nickels 104 quarters; 46 nickels 95 quarters; 55 nickels

OpenStudy (anonymous):

104 quarters and 46 nickels . 104*.25=26$ 46*.05=2.30$ $26+$2.30=28.30

OpenStudy (anonymous):

a quarter is 0.25 and nickel is 0.05 so just multiple and add the answer to check

OpenStudy (anonymous):

Amy mixes 12 liters of 16% acid solution with a 40% acid solution which results in 32% acid solution. If x represents the number of liters of 40% solution used, which equation can be used to find the number of liters of 40% acid solution in the mixture? (Points : 1) 1.92 + 0.32x = 0.4(x + 12) 3.84 + 0.4x = 0.16(x + 12) 4.8 + 0.32x = 0.4(x + 12) 1.92 + 0.4x = 0.32(x + 12)

OpenStudy (anonymous):

I think its d

OpenStudy (anonymous):

d

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Amy mixes 12 liters of 16% acid solution with a 40% acid solution which results in 32% acid solution. Find the number of liters of 40% solution in the new mixture. (Points : 1) 36 liters 24 liters 16 liters 4.8 liters

OpenStudy (anonymous):

0.16(12) + 0.4x = 0.32(12 + x) Solve for x

OpenStudy (anonymous):

0.16(12) + 0.4x = 0.32(12 + x) 1.92 + 0.4x = 3.84 + 0.32x -1.92 -1.92 0.4x = 1.92 + 0.32x -0.32 -0.32 0.08x = 1.92 ---------- 0.08 0.08 x = 24 liters

OpenStudy (anonymous):

thanks